Debt payments of $1,100 and $2,700 are due i. five months and ten months , respectively. what single layment is required to settle both debts in one month? Assume a simple interest rate of 5.60% p.a and use one month from now as the focal date.
Answer #1

Let's calculate the present value

Debt payment in five months = $1100

Debt payment in ten months = $2700

a = 5-1 = 4 months

b = 10-1 = 9 months

r = monthly interest rate

= 5.6 ÷ 12 = 0.46666 %

Present value =( Debt payment in 5 months ÷ [ 1+r ] ^ a ) + ( Debt payment in 10 months ÷ [ 1+r ] ^ b )

= ( $1100 ÷ [ 1+0.46666% ] ^ 4 ) + ( $ 2700 ÷ [ 1+0.46666% ] ^ 9 )

= ( $1100 ÷ 1.01879775 ) + ( $2700 ÷ 1.04279263 )

= $ 1079.70399 + $ 2589.20127

= $ 3668.90526

Single payment required = $ 3669

Please provide feedback if it was helpful.